OpenStudy (anonymous):

4 8/10 divided by 4/5

OpenStudy (anonymous):

first convert the mixed number into an improper fraction by multiplying the whole number by the denominator then adding the numerator: ie, (4*10)+8 = 40 +8=48 so we have 48/10 divided by 4/5 now when you dividing fractions, you have to invert the divider and multiply the inverted form by the other fraction...thus we have \[(48/10)*(5/4)\](since 4/5 is the divider) bow it is time to reduce as much as possible... (48/10)*(5/4) can be written as: (48/4)*(5/10)= 12*(1/2)=12/2=6... thus your answer is 6...